OpenStudy (anonymous):

sec θ =5 , find sin (2θ)

OpenStudy (anonymous):

sin2x=2(sinx)(cosx) secx=5 1/cosx =5 cosx=1/5  x can be in quadrant 1 or 4 for cosx is +ve for quadrant 1, both sin and cos is +ve sinx=(sqrt24)/5 cosx=1/5 sin2x=2sinxcosx=2((sqrt24)/5)(1/5) for quadrant 4, sinx is -ve. sin2x=2(sinx)(cosx)=2(-(sqrt24)/5)(1/5)